# Mammo-Clustering:A Weakly Supervised Multi-view Global-Local Context Clustering Network for Detection and Classification in Mammography

### v1

@@ IV Discussion
Figure 5: From figure a to d, the left half of each image shows the original mammogram with annotated suspicious lesions, while the right half presents our contextual clustering visualization, akin to a CNN heatmap and a VIT attention map, with the suspicious lesion locations also outlined. This figure clearly shows that our Context Clustering approach effectively identifies and groups suspicious lesion areas in mammography.
Figure 5 reveals that lesions in mammography predominantly appear in clustered forms. The Coc feature extraction paradigm utilizes clustering of point sets based on color and location information, making it adept at identifying the shape and position of such lesions. This paradigm offers simplicity, resulting in excellent interpretability and generalizability. Therefore, we believe that the Coc network has a strong capability to learn the prior structure of lesions in mammography images.
Single-View GMIC Multi-View GMIC Ours
MDR 0.291 0.277 0.177
TABLE V: The performance of different models in terms of missed detection rate.
We introduce an additional evaluation metric, the missed detection rate (MDR):
$\textit{MDR}=\frac{N_{miss}}{N_{gt}}$
MDR is defined as the percentage of the number of unrecognized suspicious lesion areas $N_{miss}$ relative to the total number of suspicious lesion areas $N_{gt}$. Because, in clinical practice, we are more concerned about lesions being undetected, i.e., false negatives, rather than false positives.
From Table V, we can see that our model achieved the lowest missed detection rate of 0.177, surpassing other model with rates close to 0.1, demonstrating its potential. The other two models, like ours, are weakly supervised methods capable of identifying lesion locations using only classification labels, making them suitable for comparison.
@@ V Conclusions
In this study, we developed a novel weakly supervised multi-view model for early breast cancer screening using mammography images. Unlike conventional feature extraction paradigms such as CNNs and Transformers, our approach employs a context clustering-based method. This strategy allows for the integration of feature-based local information with patch-based local information, enhancing the model’s ability to capture nuanced image details. Furthermore, by incorporating multi-view mammography image features, our model effectively leverages complementary information from different perspectives. This comprehensive approach addresses the limitations of single-view analysis and improves diagnostic accuracy. The model’s performance was rigorously evaluated on two publicly available datasets, Vindr-Mammo and CBIS-DDSM, where it achieved state-of-the-art accuracy, demonstrating its potential as a robust tool for early breast cancer detection.

### v2

@@ IV Discussion
Although our model achieves state-of-the-art performance on two public datasets, the varying performance of MIL across different datasets has drawn our attention. We believe the poor performance of MIL on the Vindr-Mammo dataset is due to its long-tail distribution. Addressing the impact of long-tail distribution and exploring MIL’s broader performance on well-balanced datasets will be an interesting research direction in the future.
Exploring ways to better assist clinicians in screening tasks to reduce their workload also represents a more promising research avenue.
@@ V Conclusions
This study presents a novel, weakly supervised multi-view tri-level information fusion framework for early breast cancer screening using mammography images. Unlike conventional feature extraction paradigms such as CNN and ViT, our approach leverages a context-clustering-based methodology. This paradigm enhances computational efficiency and facilitates easier association of structural or pathological features, making it suitable for clinical tasks in mammography. Additionally, through the tri-level information fusion framework, the proposed model effectively integrates complementary information from different levels, significantly enhancing diagnostic accuracy. Comprehensive evaluations conducted on two publicly available datasets, Vindr-Mammo and CBIS-DDSM, showcase the model’s exceptional performance, achieving state-of-the-art accuracy. These results underscore the potential of the proposed framework as a robust and scalable solution for early breast cancer detection, offering significant promise for deployment in clinical settings.